OpenStudy (anonymous):

3.04]What is the Least Common Denominator (LCD) of nine-sixteenths and nineteen twenty-fourths?

OpenStudy (mathmale):

Examine the denominators, 16 and 24. Factor each and look for the common factor or factors. 16 = 2*8; 24 = 3*8. From this info, I multiply 16 by 3 and 24 by 2, obtaining 48 in both cases. The LCD is thus 48. Note that you can divide 48 by 16 and obtain 3; you can divide 48 by 24 and obtain 2.